OpenStudy (anonymous):

I need help .-. How do you substitute x=y+ 7 and y-8 = 2x ? Thank you.

OpenStudy (anonymous):

(It's solving systems btw)

OpenStudy (anonymous):

ello love :D need some help? :)

OpenStudy (anonymous):

Yes please . :)))

OpenStudy (anonymous):

alright so we have the value of x which is: x=y+7 so we begin by substituting that equation for x in the other equation like so: y-8=2(y+7) follow me?

OpenStudy (anonymous):

So we distribute the equation right?

OpenStudy (anonymous):

yes we distribute the 2 into the parenthesis :)

OpenStudy (anonymous):

 yes?